OpenStudy (he66666):

Vectors: Determine the value for x such that the points A(-1, 3, 4), B(-2, 3, -1), and C(-5, 6, x) all lie on a plane that contains the origin. answer: x=-7 How do you solve this? Especially since it says it should contain the origin..

OpenStudy (anonymous):

There is some equation in three variables that they all fit: Aw + By + Cz = D Since it includes the origin, then D = 0

OpenStudy (anonymous):

sorry, A, B, and C have nothing to do with the point names in your problem statement!

OpenStudy (anonymous):

-1w +3y +4z = 0 -2w + 3y -1z = 0 -5w +6y -xz = 0 hmm... four vars and three eq's...

OpenStudy (he66666):

I don't think our teacher does it think way. She uses linear combinations to solve these questions :S so (-5, 6, x) = a(-1, 3, 4) + b(-2, 3, -1)

OpenStudy (he66666):

this*